# Efficient Reinforcement Learning for Optimal Control with Natural Images

### v1

@@ 3.2 Benchmark Performance with Length of Horizon
The optimal and greedy policies given in Tables 1 and 2 turn out to be stationary policies for all values of $p$. This is confirmed using an infinite horizon analysis. The cost of each policy is found in the Appendix by introducing a discount factor and applying policy evaluation in the infinite horizon limit ($N\rightarrow\infty$). These calculations are in complete agreement with value iteration applied to the discounted problem (not shown). In addition, the ratio of costs of the greedy and optimal policies in Tables 1 and 2 is given in the Appendix for the case of an infinite horizon, and the result is completely consistent with our discussion of these policies for finite horizons. For example, in the infinite horizon limit the expected total cost ratio of greedy and optimal policies becomes 2 when $p=0$, 5 when $p=0.75$, and diverges when $p=1$. For finite horizons, the first result ($p=0$) is most easily seen in Fig 3 when $N=30$; giving a ratio of $20/10=2$. The final result ($p=1$) is also seen in Fig 3 when $N=30$, giving a ratio of $29/1=29$; which will clearly diverge as $N\rightarrow\infty$. The result for $p=0.75$ can be seen from the ratio of expected total costs of the greedy and optimal results in Fig 7.
The optimal and greedy policies undergo more complicated dynamics over shorter horizons when different initial states are chosen. For an initial state given by $((0,1),(0,0))$ (i.e., where the tracker starts close to the target) the bottom pair of curves in Fig. 4 reach the stationary optimal and greedy policies in Tables 1 and 2 at very short horizons. As the initial distance between the tracker and target increases for positive coordinates (middle and top pairs of curves in Fig. 4), the optimal and greedy policies take longer and longer to reach these stationary policies. For example, the top pair of curves has not reached their stationary policies until the horizon is around 14 time periods. These results suggest the dynamics of all initial states with one positive coordinate and one non-negative coordinate eventually reach the stationary policies in Tables 1 and 2 provided the horizon is long enough.
@@
In Fig. 4, the greedy policy is shown to be an optimal policy when the initial state of the tracker is far from the target. Reaching a stationary policy only happens when the tracker gets close enough to the target (within a distance of one). Once the tracker is close to the target the greedy policy is no longer optimal (as shown for longer horizons in Fig. 4). For initial states where both coordinates are negative, the tracker is always behind the target and can never catch up. A simple greedy approach is then sufficient to follow the target, ensuring that the greedy policy is an optimal policy in this case. A more detailed discussion of initial state dependence is given in Sec. 3.5.
@@ 3.3 Efficient Image Representations for Reinforcement Learning
To find efficient representations for natural image sequences used in reinforcement learning, image patches from the benchmark were transformed into overcomplete sparse codes and whitened complete codes. A set of 48 natural images was taken from the database “Natural Scene Statistics in Vision Science” by Geisler and Perry (2011). Each image was cropped to $2844\times 2844$ pixels to make it square, then converted to grayscale and double precision using the Matlab functions rgb2gray and im2double, and subsequently discretized into square image patches of side-length $a$ (in pixels). The number of image patches in a single image was $\lfloor 2844/a\rfloor^{2}$; where the function $\lfloor$.$\rfloor$ . ⌋ rounds down to the nearest integer.
Overcomplete sparse codes were constructed using the method described in Loxley (2021) and summarized in the Appendix. This method relies on the two-dimensional Gabor function, originally used to model simple-cell receptive field profiles in the primary visual cortex (Hubel and Wiesel, 1959; Jones and Palmer, 1987; Daugman, 1985, 1989). The Gabor function parameters are chosen by drawing samples from a tractable multivariate probability distribution that well-approximates natural image statistics (Loxley, 2017). A sparse code is given by the hidden-layer activities of an autoencoder with an input/output that comprises the same image patch, and connection weights given by 2D Gabor functions adapted to natural image statistics. The encoding requires solving a linear least-squares problem, while decoding is a simple matrix-vector multiplication. The size of the sparse code depends on the number of Gabor functions used to generate the code. To generate a ×$\times 64$ overcomplete sparse code, the number of Gabor functions required is $64\times a^{2}$ for an image patch of $a^{2}$ pixels. For reasons that will become apparent, we would like to choose $a$ to maximize the number of image patches available in a $2844\times 2844$ image, whilst ensuring the size of a ×$\times 64$ overcomplete sparse code is larger than this number; i.e., $64\times a^{2}>\lfloor 2844/a\rfloor^{2}$. The unique solution to this problem is $a=19$, leading to 22201 image patches of 361 pixels each.
@@ 3.4 Benchmark Performance with Number of States
Training a neural network on efficient image representations takes fewer computational resources than training it directly on images (Loxley, 2021). This is shown in Figs. 5 and 6, where the number of least-squares iterations required to store a given number of cost-to-go values in a linear network is shown for network inputs given by image patches and image-patch representations. Each state of the benchmark corresponds to a unique image patch and its associated cost-to-go value. By adjusting the network weights during training, this association is stored in the network in order to determine an optimal policy. In a linear network, the number of network weights determines the storage capacity of the network; that is, the maximum number of cost-to-go values it can store. In turn, the number of weights is equal to the number of pixel inputs to the network; since each input is multiplied by its corresponding weight and summed together to yield the network output. Therefore, the storage capacity of a network can be increased by increasing the number of pixel inputs and network weights. One way to do this for a fixed-sized image patch is to generate an overcomplete representation of the image patch, and use this as input to the network instead. This approach will be used to solve the benchmark tracking task as the number of states, and therefore, the number of stored cost-to-go values, increases. Network training is carried out using the Matlab lsqr function running on an Nvidia GPU, with the number of least-squares iterations until convergence (within $1\times 10^{-6}$ of the least-squares objective), averaged over 48 independent trials, shown along the vertical axis.
@@
Figure 5: Number of least-squares iterations to store a number of cost-to-go values in a linear network using different image-patch representations (see legend). The least-squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 400 pixels.
@@
In Fig. 5, the size of each image patch was chosen to be 400 pixels instead of 361 pixels for visualization purposes when the number of stored cost-to-goes is relatively small. When the input is either an image patch or a complete representation of an image patch, a linear network will have exactly 400 inputs and 400 weight parameters, allowing it to store a maximum of 400 cost-to-go values. The curve with black circles (×$\times 1$ raw image), and the curve with blue circles (×$\times 1$ whitened image) show each of these cases, and both curves get close to 400 stored cost-to-goes. However, input given by a whitened image generally takes far fewer least-squares iterations to store the same number of cost-to-goes as input given by a raw image, making it an efficient image representation. The curve with black triangles (×$\times 4$ raw image) shows results for a natural image patch that has been upscaled by a factor of four using bicubic interpolation via Matlab’s imresize function, while the curve with blue triangles (×$\times 4$ sparse code) shows results for an overcomplete sparse code. Each of these representations has 1600 pixels (i.e., $4\times 400$ pixels, treating Gabor coefficients in a sparse code the same as double-precision pixels in an image). Increasing the number of weights to 1600 allows a linear network to store a maximum of 1600 cost-to-go values. However, the best result for the ×$\times 4$ resized image remains below 400 stored cost-to-goes, and uses approximately the same number of least-squares iterations as the best result for the ×$\times 4$ sparse code; which gets close to 1600 stored cost-to-goes (the best result in Fig. 5 is at 1323 cost-to-goes for 2701 least-squares iterations). These results confirm that an efficient image representation must be both overcomplete and sparse in order to increase the storage capacity of a linear network. The key role of sparsity is discussed in Loxley (2021), where it is shown how a sparse code that decorrelates neighbouring pixels can lead to an increase in the column-space dimension of the design matrix for the least-squares problem.
@@
Figure 6: Number of least-squares iterations to store a number of cost-to-go values in a linear network using different image-patch representations (see legend). The least-squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 361 pixels.
@@
Surprisingly, if we continue to increase the overcompleteness of image representations, the network storage capacity continues to increase. This is shown in Fig 6 for network inputs given by ×$\times 4$, ×$\times 16$, and ×$\times 64$ overcomplete sparse codes. The size of each image patch is 361 pixels; so the ×$\times 4$ sparse code has $4\times 361=1444$ pixels, the ×$\times 16$ sparse code has $16\times 361=5776$ pixels, and the ×$\times 64$ sparse code has $64\times 361=23104$ pixels. These quantities are shown as dashed vertical lines in Fig 6. The general result in Fig 6 is that the number of stored cost-to-goes (i.e., the network storage capacity) increases as the overcompleteness of an image representation increases. The first vertical dashed line at 361 cost-to-go values corresponds to the maximum storage capacity for a linear network with inputs given by complete codes. The other vertical dashed lines correspond to maximum storage capacities for inputs given by ×4,×16$\times$ 4,\ $\times$ 16,× 4 , × 16 , and ×$\times 64$ overcomplete sparse codes, respectively.
@@
Representation ##\## LS Iterations ##\## Stored Cost-To-Goes
×$\times 1$ whitened image 177 243
×$\times 4$ sparse code 2963 1323
×$\times 16$ sparse code 4527 5043
×$\times 64$ sparse code 9679 19683
×$\times 64$ sparse code 20009 21675
Table 3: (Top) Four data points in Fig 6. (Bottom) One data point not shown in Fig  6.
Some variation in the average number of least-squares iterations can be seen in Table 3, particularly for the ×$\times 4$ sparse code. This is due to the relative proximity of each data point to its limiting case given by each of the vertical dashed lines. For example, the best result for the ×$\times 4$ sparse code is the data point at 1323 stored cost-to-go values, which is closer to its limiting value of 1444 than any other image representation is to its limiting value, and hence requires a relatively large number of least-squares iterations as a result of the larger condition number of its Hessian matrix.
Now we are in a position to look at how the benchmark performs as the number of states increases. Increasing the number of states increases the size of the region for target tracking, and the number of stored cost-to-go values required for an optimal solution. Following from the results in Fig 6, we would expect that efficient image representations should increase the number of states over which optimal target tracking is possible. Fig 7 shows this result: the expected total cost of tracking a target with $p=0.75$ over thirty time periods is shown for various image representations as the number of states of the benchmark is increased. The initial states are $((0,1),(0,0))$ for the optimal tracker, and $((0,0),(0,0))$ for the greedy tracker. For comparison, the exact expected total costs for optimal and greedy target tracking are shown as the two right columns in Fig 7.
@@
According to Fig 7, a representation given by a ×$\times 1$ whitened image can solve the benchmark optimally for 243 states, but not for 363 states – where the expected total cost exceeds that for greedy target tracking. The reason is, as in Fig 6, a complete code stores a maximum of 361 cost-to-go values – which is less than 363. Similarly, a ×$\times 4$ sparse code can solve the benchmark optimally for 1323 states, but not for 1587 ($>1444$) states; a ×$\times 16$ sparse code can solve the benchmark optimally for 5547 states, but not for 6075 ($>5776$) states; and a ×$\times 64$ sparse code can solve the benchmark optimally for 21675 states. This number was the largest number of image patches tested in this work. Values for the numbers of states shown in Fig 7 result from the parameterization of the benchmark set $D$ implemented here, as previously mentioned.
@@ 3.5 Benchmark Performance for Different Initial States
We now look at the effect of starting at different initial states in the reinforcement learning benchmark. This is done by starting the benchmark in an initial state, and then running it with optimal and greedy policies until reaching the horizon. The total cost of the optimal and greedy policies can then be compared. This is repeated for each state of the benchmark.
@@
Training the benchmark with a ×$\times 64$ sparse code leads to 21675 possible initial states. The expected total cost of optimal and greedy policies is shown in Fig 8 for each initial state leading to a suboptimal greedy policy. These initial states are indexed from 1 to 10880, comprising roughly half of the 21675 available initial states. For the remaining 10795 initial states (not shown) the greedy policy is the optimal policy. The benchmark horizon is 200 time periods, and the Markov chain parameter is $p=0.4$. The policy found using fitted value iteration exactly matches the optimal policy found using dynamic programming, as shown in Fig 8 (Inset) for initial states indexed from 4000 to 4050. Fig 9 shows the distribution of cost differences, given by the cost of the greedy policy minus the cost of the optimal policy, for each initial state in Fig 8. These cost differences increase further for longer horizons.
@@
Figure 8: Expected total cost of optimal and greedy policies for each initial state leading to a suboptimal greedy policy. These initial states are indexed from 1 to 10880. Fitted VI is the policy found using fitted value iteration. Inset shows a close-up of initial states indexed from 4000 to 4050. The benchmark horizon is 200 time periods, and $p=0.4$ (see Text for details).
Figure 9: Fitted distribution of cost differences, given by the cost of the greedy policy minus the cost of the optimal policy, for each initial state in Fig 8. The fitted distribution is a kernel density estimate using the Gaussian distribution with bandwidth = 3.47.
@@
Looking at a few specific initial states helps to understand Fig 8. The lowest cost optimal policy starts at state $((0,0),(1,1))$ and has an expected total cost of 54. Since this state is part of the stationary policy in Table 1, we can find its total cost for $p=0$; which is $(200-2)/3\times 1=66$. From Fig 3, we also know the optimal cost for $p=0.4$ would be less than that for $p=0$, justifying the order relation $54<66$. Further, making use of the infinite horizon result from the Appendix, we predict an approximate value for the expected total cost of the greedy policy as: $(1+1/(1-p))\times 54=144$ for $p=0.4$. This exactly matches the value in Fig 8.
@@
Now that we have a reasonable theoretical justification for the lowest cost policy in Fig 8, let us attempt to understand the highest cost policy. This policy starts at state $((42,-42),(1,1))$ and has an expected total cost of $185870$ for the optimal policy, and $185890$ for the greedy policy. We can understand this as follows. Each coordinate for $c_{k}-t_{k}$ in the benchmark has the range $[-42,42]$ when the number of states is 21675. Starting at state $((42,-42),(1,1))$ means the tracker is very far behind the target, but it is also very far to the right of the target. Consider the corresponding situation when $p=0$: during three time periods, the target will move right once and up twice. If the tracker simply moves up three times over the same three time periods, it can reach the target within $42\times 3=126$ time periods, and apply the stationary optimal policy thereafter. There is no other initial state with a suboptimal greedy policy that can have a larger expected total cost. The same argument holds approximately when $p=0.4$.
@@
For the remaining 10795 initial states where the greedy policy is the optimal policy, the lowest cost policy starts at state $((0,0),(0,0))$ and has an expected total cost of 145. This is close to the expected total cost of the greedy policy for the initial state $((0,0),(1,1))$ just discussed. Predictably, the highest cost policy starts with the tracker as far behind the target as possible, in state $((-42,-42),(1,1))$, and has an expected total cost of $701100$.
3.6 Benchmark Performance with Alternative Sampling Strategies

### v2

@@ 4.2 Benchmark Performance with Length of Horizon
The optimal and greedy policies given in Tables 1 and 2 turn out to be stationary policies for all values of $p$. This is confirmed using an infinite horizon analysis. The cost of each policy is found in the Appendix by introducing a discount factor and applying policy evaluation in the limit of an infinite horizon (i.e., $N\rightarrow\infty$). These calculations also agree with value iteration applied to the discounted problem (not shown). Furthermore, the cost ratio of the greedy to optimal policies in Tables 1 and 2 is given in the Appendix in the limit of an infinite horizon for all values of $p$. The result is completely consistent with our discussion of these policies for finite horizons. For example, this ratio becomes 2 when $p=0$, 5 when $p=0.75$, and diverges when $p=1$. For finite horizons, the first result ($p=0$) is most easily seen in Fig 4 when $N=30$; giving a ratio of $20/10=2$. The final result ($p=1$) is also seen in Fig 4 when $N=30$, giving a ratio of $29/1=29$. As $N$ increases, this numerator will increase, diverging in the limit as $N\rightarrow\infty$. The result for $p=0.75$ can be seen from the ratio of expected total costs of the greedy and optimal results in Fig 8.
The optimal and greedy policies lead to more complicated dynamics over shorter horizons when different initial states are chosen. For an initial state given by $((0,1),(0,0))$ (i.e., where the tracker starts close to the target) the bottom pair of curves in Fig. 5 reach the stationary optimal and greedy policies in Tables 1 and 2 at very short horizons. As the initial distance between the tracker and target increases for positive coordinates (middle and top pairs of curves in Fig. 5), the optimal and greedy policies take longer and longer to reach these stationary policies. For example, the top pair of curves has not reached their stationary policies until the horizon is around 14 time periods. These results suggest the dynamics of all initial states with one positive coordinate and one non-negative coordinate will eventually reach the stationary policies given in Tables 1 and 2 provided the horizon is long enough.
@@
In Fig. 5, the greedy policy is shown to be an optimal policy when the initial state of the tracker is far from the target. Reaching a stationary policy only happens when the tracker gets close enough to the target (i.e., within a distance of one). Once the tracker is close to the target the greedy policy is no longer optimal (as shown at longer horizons in Fig. 5). For initial states where both coordinates are negative, the tracker is always behind the target and can never catch up. A simple greedy approach is then sufficient to follow the target, ensuring the greedy policy is an optimal policy in this case. A more detailed discussion of initial state dependence is given in Sec. 4.5.
@@ 4.3 Efficient Image Representations for Reinforcement Learning
Image representations for natural image sequences used in reinforcement learning were found by transforming image patches from the benchmark into overcomplete sparse codes and whitened complete codes. A set of 48 natural images was taken from the database “Natural Scene Statistics in Vision Science” by geisler. Each image was cropped to $2844\times 2844$ pixels, then converted to grayscale and double precision using the Matlab functions rgb2gray and im2double, before being discretized into square image patches of side-length $a$ (measured in pixels). The number of image patches in a single image was therefore $\lfloor 2844/a\rfloor^{2}$; where the function $\lfloor.\rfloor$ rounds down to the nearest integer.
Overcomplete sparse codes were constructed using the sparse autoencoder described in Figure 1 and detailed in the Appendix (see also loxley2). This method relies on the two-dimensional Gabor function, originally used to model simple-cell receptive field profiles in the primary visual cortex (hubel; jones; daugman1; daugman2). The Gabor function parameters were chosen by drawing samples from a tractable multivariate probability distribution that well-approximates natural image statistics (loxley1). The size of the sparse code depends on the number of Gabor functions in the overcomplete basis used to reconstruct an image input. To generate a $\times 64$ overcomplete sparse code, the number of Gabor functions required is $64\times a^{2}$ for an image patch of $a^{2}$ pixels. In order to manage the storage capacity of our network we enforce the inequality $m>n$; where $m$ is the size of an overcomplete sparse code, and $n$ is the number of images (i.e., the number of cost-to-goes stored in the network). For this reason, we choose $a$ to maximize the number of image patches available in a $2844\times 2844$ image, whilst ensuring the size of a $\times 64$ overcomplete sparse code is larger than this number; i.e., $64\times a^{2}>\lfloor 2844/a\rfloor^{2}$. The unique solution to this problem is $a=19$, leading to 22201 image patches of 361 pixels each.
@@ 4.4 Benchmark Performance with Number of States
Training a neural network on efficient image representations takes fewer computational resources than training it directly on images (loxley2). This can be seen in Figs. 6 and 7, where the number of least squares iterations required to store a given number of cost-to-go values in a linear network is shown for network inputs given by image patches, and representations of image patches. Each state of the benchmark is given by a unique image patch that is associated with a cost-to-go value. By adjusting the network weights during training, this association is stored in the network in order to determine an optimal policy. In a linear network, the number of network weights determines the storage capacity of the network; that is, the maximum number of cost-to-go values it can store. In turn, the number of weights is equal to the number of pixel inputs to the network; since each input is multiplied by its corresponding weight and summed together to yield the network output. Therefore, the storage capacity of a network can be increased by increasing the number of pixel inputs and network weights. One way to do this for a fixed-sized image patch is to generate an overcomplete representation of an image patch, and use this as input to the network instead – as shown in Figure 1. This approach will be used to solve the benchmark tracking task as the number of states and stored cost-to-go values increases. Network training is carried out using the Matlab lsqr function running on an Nvidia GPU, with the number of least squares iterations until convergence (within $1\times 10^{-6}$ of the least squares objective), averaged over 48 independent trials, shown along the vertical axis.
@@
Figure 6: Number of least squares iterations to store a number of cost-to-go values in a linear network using different image patch representations (see legend). The least squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 400 pixels in this figure.
@@
In Fig. 6, the size of each image patch was chosen to be 400 pixels instead of 361 pixels for visualization purposes when the number of stored cost-to-goes is relatively small. When the network input is either a raw image patch or a complete representation of an image patch, a linear network will have exactly 400 inputs and 400 weight parameters, allowing it to store a maximum of 400 cost-to-go values. The curves for a $\times 1$ raw image and a $\times 1$ whitened image show these cases, and both curves get close to 400 stored cost-to-goes. However, input given by a whitened image generally takes far fewer least squares iterations to store the same number of cost-to-goes as a raw image, making it a computationally efficient image representation. The curve for a $\times 4$ raw image shows results for a natural image patch that has been upscaled by a factor of four using bicubic interpolation via Matlab’s imresize function. The curve for a $\times 4$ sparse code shows results for an overcomplete sparse code. Each of these representations has 1600 pixels (i.e., $4\times 400$ pixels, treating the coefficients of a Gabor function basis the same as double precision pixels in an image). Increasing the number of weights to 1600 allows a linear network to store a maximum of 1600 cost-to-go values. However, the best result for the $\times 4$ resized image remains below 400 stored cost-to-goes, and uses approximately the same number of least squares iterations as the best result for the $\times 4$ sparse code; which gets closer to 1600 stored cost-to-goes (the best result in Fig. 6 is 1323 cost-to-goes for 2701 least squares iterations).
These results confirm that a computationally efficient image representation must be both overcomplete and sparse in order to increase the storage capacity of a linear network. The key role of sparsity is discussed in loxley2, where it is shown how a sparse code decorrelates neighbouring pixels in an image and leads to an increase in the column-space dimension (rank) of the design matrix for the associated least squares problem. A sparse code also leads to a well-conditioned Hessian matrix, reducing the number of least squares iterations required for convergence. Since standard decorrelation techniques do not generalize to overcomplete representations, an overcomplete sparse code seems to be a minimum requirement for increasing neural network storage capacity efficiently.
@@
Figure 7: Number of least squares iterations to store a number of cost-to-go values in a linear network using different image patch representations (see legend). The least squares objective is found within a tolerance of $1\times 10^{-6}$, and the number of iterations are averaged over 48 independent trials. Each image patch is 361 pixels.
@@
Surprisingly, if we continue to increase the overcompleteness of image representations, the network storage capacity continues to increase. This is shown in Fig 7 for network inputs given by $\times 4$, $\times 16$, and $\times 64$ overcomplete sparse codes. The size of each image patch is 361 pixels; so the $\times 4$ sparse code has $4\times 361=1444$ pixels, the $\times 16$ sparse code has $16\times 361=5776$ pixels, and the $\times 64$ sparse code has $64\times 361=23104$ pixels. These quantities are shown as dashed vertical lines in Fig 7. The general result in Fig 7 is that the number of stored cost-to-goes (i.e., the network storage capacity) increases as the overcompleteness of an image representation increases. The first vertical dashed line at 361 cost-to-go values corresponds to the maximum storage capacity of a linear network with inputs given by complete codes. The other three vertical dashed lines correspond to maximum storage capacities for inputs given by $\times 4,\ \times 16,$ and $\times 64$ overcomplete sparse codes, respectively.
@@
Representation #\# LS Iterations #\# Stored Cost-To-Goes
$\times 1$ whitened image 177 243
$\times 4$ sparse code 2963 1323
$\times 16$ sparse code 4527 5043
$\times 64$ sparse code 9679 19683
$\times 64$ sparse code 20009 21675
Table 3: (Top) Four data points in Fig 7. (Bottom) One data point not shown in Fig  7.
Some variation in the average number of least-squares iterations can be seen in Table 3, particularly for the $\times 4$ sparse code. This is due to the relative proximity of each data point to its limiting case given by each of the vertical dashed lines. For example, the best result for the $\times 4$ sparse code is the data point at 1323 stored cost-to-go values, which is closer to its limiting value of 1444 than any other image representation is to its limiting value, and hence requires a relatively large number of least squares iterations as a result of the larger condition number of its Hessian matrix.
Now we are in a position to look at how the benchmark performs as the number of states increases. Increasing the number of states increases the size of the region over which a target may be tracked in the benchmark, as well as the number of stored cost-to-go values required for an optimal solution. Following from the results in Fig 7, we would expect that using efficient image representations should increase the number of states over which optimal target tracking is possible. Fig 8 shows this result: the expected total cost of tracking a target with $p=0.75$ over thirty time periods is shown for various image representations as the number of states of the benchmark is increased. The initial states are taken from Tables 1 and 2 and are $((0,1),(0,0))$ for the optimal tracker, and $((0,0),(0,0))$ for the greedy tracker. For comparison, exact results are shown for the expected total costs of optimal and greedy target tracking in the two right columns in Fig 8.
@@
According to Fig 8, a representation given by a $\times 1$ whitened image can solve the benchmark optimally for 243 states, but not for 363 states – as the expected total cost now exceeds that for greedy target tracking. The reason is, as in Fig 7, a complete code stores a maximum of 361 cost-to-go values which is less than 363. Similarly, a $\times 4$ sparse code can solve the benchmark optimally for 1323 states, but not for 1587 (>1444>1444) states; a $\times 16$ sparse code can solve the benchmark optimally for 5547 states, but not for 6075 (>5776>5776) states; and a $\times 64$ sparse code can solve the benchmark optimally for 21675 states. This was the largest number of image patches tested in this work. Actual values for the number of states shown in Fig 8 are due to the specific parameterization of the set $D$ implemented here.
@@ 4.5 Benchmark Performance for Different Initial States
We now look at the effect of starting at different initial states in the reinforcement learning benchmark. This is done by starting the benchmark in an initial state, and then running it with either an optimal or greedy policy until reaching the horizon. The total cost of each policy can then be compared. This is repeated for each state of the benchmark.
@@
Training the benchmark with a $\times 64$ sparse code leads to 21675 possible initial states. The expected total cost of optimal and greedy policies is shown in Fig 9 for each initial state that leads to a suboptimal solution under the greedy policy. These initial states are indexed from 1 to 10880 and comprise roughly half of the 21675 available initial states. For the remaining 10795 initial states (not shown) the greedy policy is the optimal policy. The benchmark horizon is 200 time periods, and the Markov chain parameter is $p=0.4$. The policy found using fitted value iteration matches the optimal policy found using exact dynamic programming, as shown in Fig 9 (Inset) for initial states indexed from 4000 to 4050.
@@
Figure 9: Expected total cost of optimal and greedy policies for each initial state leading to a suboptimal greedy policy. These initial states are indexed from 1 to 10880. “Fitted VI” is the policy found using fitted value iteration. Inset shows a close-up of initial states indexed from 4000 to 4050. The benchmark horizon is 200 time periods, and $p=0.4$ (see text for details).
@@
Looking at a few specific initial states helps to understand Fig 9. The lowest cost optimal policy in Fig 9 starts at state $((0,0),(1,1))$ and has an expected total cost of 54. We can derive a theoretical bound for this state since it is part of the stationary policy in Table 1 (see also Figure 4). For $p=0$, its total cost is given by $(200-2)/3\times 1=66$. From Fig 4, we know the optimal cost for $p=0.4$ would be less than that for $p=0$, leading to $54<66$. We can also predict a value for the lowest cost greedy policy by making use of the infinite horizon result for the total cost ratio from the appendix: (1+1/(1p))(1+1/(1-p)). The expected total cost of the greedy policy can then be found by multiplying this ratio by the expected total cost of the optimal policy: for $p=0.4$ we get $(1+1/(1-p))\times 54=144$. This exactly matches the value in Fig 9, which is too small to see in that figure.
@@
Now that we have some theoretical justification for the lowest cost policy in Fig 9, let us attempt to understand the highest cost policy in that figure. This policy starts at state ((42,42),(1,1))((42,-42),(1,1)) and has an expected total cost of $185870$ for the optimal policy, and $185890$ for the greedy policy. We can understand this as follows. Each number in the coordinate pair for $c_{k}-t_{k}\in D$ has the range [42,42][-42,42] when the number of states is 21675. Starting at state ((42,42),(1,1))((42,-42),(1,1)) means the tracker is as far behind the target as possible, but it is also as far to the right of the target as possible. Consider the corresponding situation when $p=0$: during three time periods, the target will move right once and up twice. If the tracker simply moves up three times over the same three time periods, then it has decreased the distance to the target by one every three time periods. Therefore, it can reach the target within $42\times 3=126$ time periods, and apply the stationary optimal policy thereafter. There is no other initial state with a suboptimal greedy policy that can have a larger expected total cost. The same argument holds approximately when $p=0.4$.
@@
For the other 10795 initial states where the greedy policy is the optimal policy (not shown), the lowest cost policy starts at state $((0,0),(0,0))$ and has an expected total cost of 145. This is close to the expected total cost of the greedy policy for the initial state $((0,0),(1,1))$ just discussed. Predictably, the highest cost policy starts with the tracker as far behind the target as possible, in state ((42,42),(1,1))((-42,-42),(1,1)), and has an expected total cost of $701100$.
4.6 Benchmark Performance with Partial Training Sets